# Affordable Generative Agents

### v1

@@ 1 Introduction
Constructing an agent based on large language models (LLMs) for the simulation of believable human behavior holds substantial promise (da Rocha Costa, 2019; Park et al., 2022; Xi et al., 2023a). A believable LLM agent is characterized by its ability to interact with other agents or humans, respond to environmental changes, and generate responses and behaviors that are perceived by humans as authentic, natural, and in alignment with expectations. Such agents could be utilized for diverse applications, such as simulating the potential impacts of policies (Zhang et al., 2023a; Törnberg et al., 2023) or sociological theories on humans (Gao et al., 2023), creating Non-Player Characters (NPCs) (Laird & VanLent, 2001; Gong et al., 2023) in games, and developing embodied intelligence akin to social robots (Cherakara et al., 2023) or domestic service robots (Puig et al., 2020; Yang et al., 2023b).
@@
However, while we appreciate the potential and capabilities of these agents, their cost implications also merit attention. The crux of the matter is that real-time, long-term invocation of LLMs to handle various types of agent interactions can result in substantial cost consumption. Previous studies have highlighted this issue. For instance, Park et al. (2023) construct a virtual town, aka the “Stanford Town”, to simulate social behavior, with the entire study incurring costs of thousands of dollars. Wang et al. (2023b) mimics embodied agents, fulfilling both emotional and rational needs through interaction, with simulations of 2 to 3 agents costing several dollars per hour.
The cost of believable agent interaction has hindered its use as a low-cost solution for simulating policy impacts and sociological phenomena, especially in long-term, large-scale scenarios. The cost, which increases linearly with the number of interactions, is unacceptable, particularly when LLM agents are to be used as NPCs in an open world environment.
To address the issues above, there has been a significant amount of research (Liang et al., 2020; So et al., 2021; Miao et al., 2023; Chen et al., 2023a; Zhang et al., 2023b). Nevertheless, existing methods primarily focus on specific independent tasks, solving problems through single-round QA sessions with LLMs (Chen et al., 2023a; Zhang et al., 2023b). Consequently, there lacks solutions for continuous multi-round agent interactions and the coordination infrastructure supporting low-cost implementations. To expand, the main reasons these methods are unsuitable for believable behavior simulation include the following aspects: 1) the lack of a unified and explicit evaluation system for believable behavior in open-world scenarios, making it unsuitable for task difficulty segmentation and cost-efficient model inference (Khalili et al., 2022); 2) the need for agents to consider context and multi-round interactions, requiring large amounts of relevant information for decision-making, making it unsuitable for combining multiple independent tasks into a single prompt for single-round inference (Lin et al., 2023; Mu et al., 2023); and 3) the need to consider the cost of single interactions as well as the redundancy in large-scale interactions, which requires optimization from the framework level.
@@
Therefore, in this paper, we focus on the believable behavior interaction simulation of LLM agents in open-world scenarios and propose a low-cost framework, termed as Affordable Generative Agents. For the agent-environment interactions, we propose the Lifestyle Policy to reduce the redundant costs of generating repeated responses by agents; On the other hand, for the inter-agent interactions, we propose the Social Memory to reduce the generation of repeated information in multi-round interactions. These two techniques are distilled from our insights on how a human interacts with other entities in the real world, and can be compiled easily into existing agent simulation environments. To examine the applicability of our techniques, we have conducted extensive experiments using well-known environments, including the “Stanford Town” (Park et al., 2023) and the VirtualHome (Puig et al., 2018), to demonstrate that, while achieving the same performance, the consumption of generating believable agent behaviors can be significantly reduced.
@@ 2 Related Work
This work explores the development of efficient generative behaviors leveraging LLMs. We discuss the most related works below.
@@ 2.1 LLM Agents
The development of LLM agents has sparked a myriad of applications. Yang et al. (2023a) leverages multi-turn dialogues with agents to autonomously address NLP tasks. Deng et al. (2023), Nakano et al. (2021), and Yao et al. (2022a) integrate pre-defined APIs with LLMs to enable agents to navigate web pages effectively. Xu et al. (2023) and Light et al. (2023) deploy LLM agents in text game environments, showing the capacity to engage in complex, strategic gameplay. The most prominent among these works is the construction of LLM agents that interact in open environments to create believable behavior. For instance, Park et al. (2023) builds a virtual town simulating the social life of 25 agents, while Brohan et al. (2023) utilizes LLM to create a robot capable of completing household chores through natural language interaction. Wang et al. (2023a) implements an agent capable of playing Minecraft expertly. Besides, researchers have been focusing on enhancing its performance. Efforts have been made to optimize various modules such as reasoning (Wei et al., 2022; Wang et al., 2022; Xi et al., 2023b), planning (Song et al., 2023; Yao et al., 2022b), reflection (Shinn et al., 2023), and memory (Zhong et al., 2023; Huang et al., 2023b), or to create more realistic interaction environments (Hong et al., 2023b; Huang et al., 2023a). However, the development of affordable and believable agent interactions is missing in the literature. In this paper, we focus on achieving believable agents with the same performance at low cost.
@@ 2.2 The Cost of LLM Interaction
The cost of using LLM is directly proportional to the number of LLM invocations, the length of the prompt, and the model’s performance. The cost issue has always been a key focus in the LLM-related field. Numerous methods have been proposed to reduce the cost of using LLM, such as Cascaded Invocation of LLM (Yue et al., 2023; Chen et al., 2023a; Nottingham et al., 2023), summary (Laban et al., 2023; Arefeen et al., 2023; Mu et al., 2023), and batch (Lin et al., 2023). These methods primarily optimize the cost of single invocation of LLM for QA tasks. Zhang et al. (2023b) has built a cost-efficient multi-agent framework, but it still solves QA tasks on fixed datasets. The simulation of believable behavior in the open world requires uninterrupted calls to the LLM for interaction. Moreover, multiple modules need to combine a large amount of contextual information for LLM reasoning and planning, making cost a more prominent issue. Simply optimizing for single QA invocations cannot effectively reduce costs. Furthermore, the open world does not have a fixed evaluation method, and the difficulty of tasks is hard to categorize, rendering many existing methods unsuitable (Chen et al., 2023a; Zhang et al., 2023b). Kaiya et al. (2023) proposes generative agents with low-cost interactions, but their effectiveness is highly dependent on the tailored scenarios. Hence, there is a pressing need for a low-cost solution that can effectively handle the challenges of believable interaction simulation in various open environments using LLMs. This paper aims to address this gap by proposing a cost-efficient scheme for simulating believable behavior with LLMs in open environments.
@@ 3 Method
Figure 1: The AGA framework, including (a) Agent-environment interactions and (b) Inter-agent interactions.
To implement the AGA framework, we abstract the interaction patterns of the LLM agent into two types: agent-environment interactions and inter-agent interactions. Figure 1 provides an overview of the AGA framework. In the following sub-sections, we will describe the optimization strategies for these two categories of interactions in detail.
@@ 3.1 Agent-environment interactions
To combine the advantages of these two methods, we propose a new method – Lifestyle Policy, which aims to retain the flexible interaction capabilities of the LLMs while achieving low-cost deployment. The functionality of Lifestyle Policy is to minimize costs by reusing the similar inference processes of LLM. To this end, we devise agent-environment interactions containing two stages: Plan Decomposition and Policy Reuse, expanded as follows.
@@
The stage occurs when the agent encounters situations it has never seen before. Due to their inherent zero-shot capabilities (Kojima et al., 2022), LLM agents can respond to unseen tasks in a believable way. Given the description of the environment and agent’s information, a rough plan or task is generated by LLM. The Plan Decomposition breaks the plan down into sub-plans and then converts sub-plans into specific actions, just like other LLM multi-agent framework (Chen et al., 2023b; Li et al., 2023; Hong et al., 2023a). An additional step involves generating an executable condition based on the action sequence. The determination of execution condition is highly correlated with the environment. For instance, in the case of a domestic robot (Brohan et al., 2023), the condition of a task include the existence of the related interactive items and the consistency of their states. If the plan can be successfully executed, then the whole graph from plan to actions and corresponding condition will be updated to Lifestyle Policy.
@@
The stage occurs when the agent encounters previously seen situations. Given a plan, we first search the Lifestyle Policy for any existing plans that match or closely resemble the new one, based on cosine similarity of their embedding features. When the similarity between two plans exceeds a certain threshold, we consider them to have same meaning, albeit with minor differences in their descriptions. Subsequently, we assess whether the current environment meets the execution condition associated with the retrieved plan. If the condition is met, we proceed with the actions specified in that plan. It is important to note that a single plan may have multiple conditions tailored for different scenarios. We iterate through all the conditions until we find one that the current environment satisfies.
Figure 1(a) illustrates the complete interaction process. Upon receiving an observation that includes environmental data and the agent’s personal profile, the LLM agent formulates a plan. Initially, the Lifestyle Policy retrieves a plan with a similar description and condition met by current environment. If a match is found, the agent executes the corresponding actions. Otherwise, the LLM is invoked to decompose the plan. The LLM plan decomposition ensures that the actions generated are of high quality, believable, and self-consistent. Meanwhile, the Lifestyle Policy aims to reuse existing plans to avoid unnecessary costs.
In the lifelong simulation of LLM agents, there are a large number of redundant, repetitive, and trivial decisions, such as in Generative agents (Park et al., 2023), where each agent repeatedly makes decisions like brushing teeth, washing face, etc. Replacing the process of invoking LLM reasoning with corresponding policies can save lots of cost, which will be shown in our experiments.
@@ 3.2 Inter-agent interactions
Drawing inspiration from existing text compression technologies (Liu et al., 2023), we introduce the Social Memory module, aimed at reducing the cost of inter-agent interactions and enabling the agents to produce more appropriate dialogues. This module is comprised of three key components: Relationship, Feeling, and Summary Events. Together, these components model the characteristics involved in agent interactions. The relationship and feeling components are responsible for recording the social connections and personal impressions between agents. These records shape how the agent interacts with others, influencing their behavior and approach. they can be a word or a phrase, like “Relationship: acquaintances, Feeling: friendly”. Notably, the two components are updated after each interaction to reflect any potential changes in relationships and emotions. The summary events component, on the other hand, focuses on summarizing all events related to the interlocutors and the topics of discussion. These events are retrieved from agent’s memory, enabling agents to quickly access necessary information from condensed text, thereby enhancing interaction efficiency.
Figure 1(b) presents the procedure of a dialogue interaction. Given the previous round of conversation, the agent computes the dialogue’s embedding features and retrieves relevant events from memory. After filtering out duplicates, these events are consolidated into summary events. As more interactions occur, additional relevant events will be integrated into the summary events module, which aims to compress information densely within a constrained text area, improving the efficiency and effectiveness of information delivery. The dialogue prompt is then formed by merging information from the last round of conversation and Social Memory, and the LLM is invoked generates the next round of dialogue. Upon the conclusion of the conversation, the relationship and feeling are updated based on the content of the dialogue.
Social Memory reduces costs of conversation by compressing auxiliary information, offering precise descriptions of social information to assist agents in responding consistently and appropriately. It also supplies high-information events to aid agents in generating accurate replies relevant to the conversation. Furthermore, as illustrated in Figure 2, Social Memory updates after each interaction, providing an explicit depiction of interlocutor relationships. This mechanism can be employed to investigate the dynamics of relationships between agents, thereby analyzing the evolution of community relations and the emergence of social behaviors.
@@ 4.1 Environment
  • Generative Agents. Park et al. (2023) creates a simulated town, populating it with multiple agents powered by LLMs that interact in natural language, generating believable individual and emergent social behaviors. Agents are required to create daily schedules, which are refined into hourly plans, and then translated into minute-level activities. All interactions are executed through natural language descriptions, accompanied by corresponding emojis for representation.
@@
    We implement a LLM agent framework like Generative Agents on VirtualHome to validate our method.
@@ 4.2 Language Models Adopted
For a fair comparison, all agents are empowered with GPT-3.5-Turbo (Wu et al., 2023), same as the Generative Agents work (Park et al., 2023). In the evaluation phase, we use GPT-4 (Achiam et al., 2023) as an evaluator to assess the agent’s believable behavior.

### v2

@@ 1 Introduction
Intelligent agents achieving human-level performance in complex tasks has attracted research efforts for years Mnih et al. (2015); Ye et al. (2020a; b); Zhao et al. (2023). Recently, constructing an agent based on large language models (LLMs) for the simulation of believable human behavior has shown substantial promise (da Rocha Costa, 2019; Park et al., 2022; Lan et al., 2023; Xi et al., 2023a). A believable LLM agent is characterized by its ability to interact with other agents or humans, respond to environmental changes, and generate responses and behaviors that are perceived by humans as authentic, natural, and in alignment with expectations. Such agents could be utilized for diverse applications, such as simulating the potential impacts of policies (Zhang et al., 2023a; Törnberg et al., 2023) or sociological theories on humans (Gao et al., 2023), creating Non-Player Characters (NPCs) (Laird & VanLent, 2001; Gong et al., 2023) in games, and developing embodied intelligence akin to social robots (Cherakara et al., 2023) or domestic service robots (Puig et al., 2020; Yang et al., 2023b).
@@
However, while we appreciate the potential and capabilities of these agents, their cost implications also merit attention. The crux of the matter is that frequent and prolonged use of LLMs to handle various types of agent interactions can result in substantial cost consumption. Previous studies have highlighted this issue. For instance, Park et al. (2023) construct a virtual town, aka the “Stanford Town”, to simulate agents spontaneously exploring and interacting with the environment to produce believable behaviors in an open world, with the entire study incurring costs of thousands of dollars. Wang et al. (2023b) mimics embodied agents, fulfilling both emotional and rational needs through interaction, with simulations of 2 to 3 agents costing several dollars per hour. The cost of believable agent interaction has hindered its application in scenarios requiring large-scale usage. For example, simulating long-term, large-scale policy impacts and sociological phenomena, as well as using LLM agents as NPCs in games.
Many methods have been developed to reduce the cost of invoking LLMs, but most of the research has focused on fixed datasets, such as COQA (Reddy et al., 2019), QQP (Wang et al., 2017), and RTE (Poliak, 2020) datasets. These tasks have clear answers, well-defined difficulty levels, and independence between tasks, typically solvable through single-turn QA interactions. However, the situation is quite different for believable agents, which lack clear answers and involve interactions that require the agent’s long-term memory and coordination between different modules. These characteristics make it challenging to apply most existing methods in this context.
@@
Therefore, in this paper, we focus on the believable behavior interaction simulation of LLM agents in open-world scenarios and propose a low-cost framework, termed as Affordable Generative Agents. We abstract the behaviors of believable agents into two modes: agent-environment interactions and inter-agent interactions. For the agent-environment interactions, we propose the Lifestyle Policy to reduce the redundant costs of generating repeated responses by agents; On the other hand, for the inter-agent interactions, we propose the Social Memory to reduce the generation of repeated information in multi-round interactions. To examine the applicability of our techniques, we have conducted extensive experiments using well-known environments, including the “Stanford Town” (Park et al., 2023) and the VirtualHome (Puig et al., 2018), to demonstrate that, while achieving the same performance, the consumption of generating believable agent behaviors can be significantly reduced.
@@ 2 Related Work
This work explores the development of efficient generative behavior leveraging LLM. We discuss the most related works below.
@@ 2.1 LLM Agents
The development of LLM agents has sparked a myriad of applications. Yang et al. (2023a) leverage multi-turn dialogues with agents to autonomously address NLP tasks. Li et al. (2024) find that the performance of LLMs scales with the number of agents instantiated across a range of tasks. Deng et al. (2023), Nakano et al. (2021), and Yao et al. (2022a) integrate pre-defined APIs with LLMs to enable agents to navigate web pages effectively. Xu et al. (2023) and Light et al. (2023) deploy LLM agents in text game environments, showing the capacity to engage in complex, strategic gameplay. The most prominent among these works is the construction of LLM agents that interact in open environments to create believable behavior. For instance, Park et al. (2023) builds a virtual town simulating the social life of 25 agents, while Brohan et al. (2023) utilizes an LLM to create a robot capable of completing household chores through natural language interaction. Wang et al. (2023a) implements an agent capable of playing Minecraft expertly, when compared to traditional game-playing methods Kanervisto et al. (2022); Lin et al. (2022). Efforts have been made to optimize various modules such as reasoning (Wei et al., 2022; Wang et al., 2022; Xi et al., 2023b), planning (Song et al., 2023; Yao et al., 2022b), reflection (Shinn et al., 2023), and memory (Zhong et al., 2023; Huang et al., 2023b), or to create more realistic interaction environments (Hong et al., 2023b; Huang et al., 2023a). However, the development of affordable and believable agent interactions is missing in the literature. In this paper, we focus on achieving believable agents with the same performance at low cost.
@@ 2.2 The Cost of LLM Interaction
The cost of using LLM is directly proportional to the number of LLM invocations, the length of the prompt, and the model’s performance. The cost issue has always been a key focus in the LLM-related field. Numerous methods have been proposed to reduce the cost of using LLM, such as Cascaded Invocation of LLM (Yue et al., 2023; Chen et al., 2023a; Nottingham et al., 2023), summary (Laban et al., 2023; Arefeen et al., 2023; Mu et al., 2023), and batch (Lin et al., 2023). These methods primarily optimize the cost of single invocation of LLM for QA tasks. Zhang et al. (2023b) has built a cost-efficient multi-agent framework, but it only solves QA tasks on fixed datasets. The simulation of believable behavior in the open world requires uninterrupted calls to the LLM for interaction. Moreover, multiple modules need to combine a large amount of contextual information for LLM reasoning and planning, making cost a more prominent issue. Simply optimizing for single QA invocations cannot effectively reduce costs. Furthermore, the open world does not have a fixed evaluation method, and the difficulty of tasks is hard to categorize, rendering many existing methods unsuitable (Chen et al., 2023a; Zhang et al., 2023b). Kaiya et al. (2023) proposes generative agents with low-cost interactions, but their effectiveness is highly dependent on the tailored scenarios. Hence, there is a pressing need for a low-cost solution that can effectively handle the challenges of believable interaction simulation in various open environments using LLMs. This paper aims to address this gap by proposing a cost-efficient scheme for simulating believable behavior with LLMs in open environments.
@@ 3 Method
Figure 1: Our proposed method for optimizing (a) Agent-environment interactions and (b) Inter-agent interactions. The gray arrows represent the baseline (Generative Agents) implementation, while the blue arrows represent our approach.
Our work is inspired by Park et al. (2023), who proposed Generative Agents, a framework for simulating believable behaviors of LLM agents in a virtual town. This framework enables agents to perceive their environment and maintain a memory stream to record past interactions and self-reflections. By retrieving relevant content from this memory stream, the framework helps agents generate believable responses. Inspired by this, We achieve generative agents in a low-cost way by optimizing agent-environment interactions and inter-agent interactions. Figure 1 illustrates our method in these two types of interactions, with gray arrows representing the Generative Agents framework’s implementation and blue arrows representing our approach. In the following sub-sections, we will describe the optimization strategies for these two categories of interactions in detail.
@@ 3.1 Agent-environment interactions
Drawing inspiration from the Case-Based Reasoning (CBR) method (Slade, 1991; Spalzzi, 2001), which addresses new problems by adapting solutions to previously solved, similar issues, we introduce a new approach – Lifestyle Policy, which aims to retain the flexible interaction capabilities of the LLMs while achieving low-cost deployment. The functionality of Lifestyle Policy is to minimize costs by reusing the similar inference processes of an LLM. To this end, we devise agent-environment interactions containing two stages: Plan Decomposition and Policy Reuse, expanded as follows.
@@
The stage occurs when the agent encounters situations it has never seen before. Due to their inherent zero-shot capabilities (Kojima et al., 2022), LLM agents can respond to unseen tasks in a believable way. Given the description of the environment and agent’s information, a rough plan or task is generated by an LLM. The Plan Decomposition breaks the plan down into sub-plans and then converts sub-plans into specific actions, just like other LLM multi-agent frameworks (Chen et al., 2023b; Li et al., 2023; Hong et al., 2023a). An additional step involves generating an executable condition based on the action sequence. The determination of execution condition is highly correlated with the environment. For instance, in the case of a domestic robot (Brohan et al., 2023), the condition of a task include the existence of the related interactive items and the consistency of their states. If the plan can be successfully executed, then the whole graph from plan to actions and corresponding condition will be updated to Lifestyle Policy.
@@
The stage occurs when the agent encounters previously seen situations. Given a plan, we use embedding models to extract features from plan descriptions and compare them with stored plan features in the Lifestyle Policy using cosine similarity. When the cosine similarity exceeds a set threshold, we consider it a match for the same plan. Setting the threshold too low may result in mismatched plans, causing agents to perform inappropriate actions, while setting it too high may lead to redundant plans being recorded in the policy. In our implementation, we set the threshold at 0.97. More details and examples of different similarity levels are provided in the Appendix A. Subsequently, we assess whether the current environment meets the execution condition associated with the retrieved plan by comparing if the interactive items in the current environment exist and have consistent states as specified in the condition. If the condition is met, we proceed with the actions specified in that plan. It is important to note that a single plan may have multiple conditions tailored for different scenarios. We iterate through all the conditions until we find one that the current environment satisfies.
Figure 1(a) illustrates the complete interaction process. Upon receiving an observation that includes environmental data and the agent’s personal profile, the LLM agent formulates a plan. For the baseline, all plans will invoke the LLM to decompose and generate corresponding actions. In our method, the Lifestyle Policy retrieves a plan with a similar description and condition met by current environment. If a match is found, the agent executes the corresponding actions. Otherwise, the LLM is invoked to decompose the plan. Due to the superiority of LLMs, the plan decomposition ensures that the actions generated are of high quality, believable, and self-consistent. Meanwhile, the Lifestyle Policy aims to reuse existing plans to avoid unnecessary costs. Due to the high similarity threshold, the replacement policy does not affect the agent’s original performance. For instance, the agent Klaus in virtual town generates an activity like "Klaus is waking up and completing his morning routine." In the Lifestyle Policy, this activity matches a similar one, "Klaus is waking up and getting ready for the day," and executes the corresponding sub-activities such as "stretching and getting out of bed," "washing his face and brushing his teeth," and "heading out the door to go to the library."
In the lifelong simulation of LLM agents, there are a large number of redundant, repetitive, and trivial decisions, such as in Generative Agents, where each agent repeatedly makes decisions like brushing its teeth, washing its face, etc. Replacing the process of invoking LLM reasoning with corresponding policies can save significant cost, which will be shown in our experiments.
@@ 3.2 Inter-agent interactions
The gray arrows in Figure 1(b) illustrate the inter-agent interactions provided by the baseline method. Specifically, agents retrieve all relevant events from the memory stream based on the previous round of conversation. LLMs then use these relevant events along with the prior conversation as a prompt to generate the next round of dialogue. However, this approach has several issues. First, trivial and repetitive relevant events can make the prompt excessively long, sometimes exceeding the LLMs’ token limit. This not only increases computational costs, especially in multi-agent, multi-turn dialogue scenarios, but also makes it difficult for the LLMs to capture key information, leading to inappropriate responses.
Drawing inspiration from existing text compression technologies (Liu et al., 2023) and social model (McCoy et al., 2011), we introduce the Social Memory module, aimed at reducing the cost of inter-agent interactions and enabling the agents to produce more appropriate dialogues. It consists of two key components: Summary Events and Relationship & Feeling. Summary Events store summaries of relevant events for each interlocutor, while Relationship & Feeling models the social connections between the agent and the interlocutor using keywords such as "Relationship: acquaintances, Feeling: friendly." In our implementation, we first retrieve relevant events based on previous dialogue content, filter out new events, and summarize them into Summary Events. These summaries increase information density while controlling text length. We then combine the agent’s personal information, event summaries, relationships, feelings, and previous dialogues into a prompt for LLMs to generate the next round of dialogue. After the dialogue ends, we input the current agent’s relationship, feelings and the summary of multiple rounds of dialogue into the LLM to update the Relationship & Feeling, as shown in Figure 2. Implementation details of each module of Social Memory are provided in the Appendix B.
Social Memory reduces costs of conversation by compressing auxiliary information, offering precise descriptions of social information to assist agents in responding consistently and appropriately. In the baseline, each conversation requires retrieving 30 to 45 relevant events, which occupy about 2000 tokens in the prompt. In contrast, our method condenses this information to approximately 100 tokens by using Summary Events. It also supplies high-information events to aid agents in generating accurate replies relevant to the conversation. Furthermore, as illustrated in Figure 2, Social Memory updates after each interaction, providing an explicit depiction of interlocutor relationships. For example, we observe that the student Klaus enjoys discussing his studies with others and occasionally mistakes the cafe owner Isabella for a fellow student in baseline. However, with our method, which employs a clear relationship update mechanism, the agent can discern social relationships during conversations. Meanwhile, This mechanism can be employed to investigate the dynamics of relationships between agents, thereby analyzing the evolution of community relations and the emergence of social behaviors.
@@ 4.1 Environment
  • Generative Agents. Park et al. (2023) creates a simulated town, populating it with multiple agents powered by LLMs that interact in natural language, generating believable individual and emergent social behaviors. Agents are required to create daily schedules, which are refined into hourly plans, and then translated into minute-level activities. All interactions are executed through natural language descriptions, accompanied by corresponding sprites and emojis for representation.
@@
    We implemented an LLM framework on VirtualHome and applied AGA to it. The detailed implementation is presented in the Appendix E.
@@ 4.2 Language Models Adopted
For a fair comparison, all agents are empowered with GPT-3.5-Turbo (Wu et al., 2023), same as the Generative Agents work (Park et al., 2023). Besides, we also conduct tests using the more advanced GPT-4 (Achiam et al., 2023).